### v1

@@ 3.3 Gaussian process models
We apply the GP model to the Superfund dataset by setting $\rho^{2}=10$ and $\sigma^{2}=1$ and using a Matérn correlation with hyperparameters $\kappa=0.2,\phi=\frac{30}{2\sqrt{0.2}}$ and the distance $d_{ij}$ between sites $i$ and $j$ is measured in meters. Specifically, $K_{ij}=\frac{2^{1-\kappa}}{\Gamma(\kappa)}(d_{ij}/\phi)^{\kappa}B_{\kappa}(d_{ij}/\phi)$ where $B_{\kappa}$ is the modified Bessel function of the third kind of order $\kappa$.
@@ 4 An encompassing weighting framework
This section presents some of our main results. Our first main result is that the GLS estimator $\widehat{\tau}_{GLS}$ in Equation 2 is a weighting estimator. Our second main result is that $\widehat{\tau}_{GLS}$ is a minimal dispersion balancing weights estimator. In particular, the weights achieve minimal variance while exactly balancing the means of measured covariates and approximately balancing the means of a hidden set of covariates. The latter property characterizes the exact nature of the adjustments done by GLS, clarifying how they emulate key features of a hypothetical experiment and how they differ from other forms of regression adjustments. We note that the results in the following propositions apply to any error covariance matrix $\bm{\Sigma}$; however, we interpret them in the context of spatial regression where $\bm{\Sigma}=\sigma^{2}\bm{I}_{n}+\rho^{2}\bm{S}$ and $\bm{S}$ encodes spatial autocorrelation. The following proposition shows that $\widehat{\tau}_{GLS}$ admits a weighting representation, as it can be re-expressed as a difference of weighted means of the treated and control outcomes.
@@
The GLS estimator of $\tau$ in Equation 2 can be expressed as
@@
with weights $\bm{w}=(w_{1},\ldots,w_{n})$ that admit the following closed-form representation
@@
Proposition 4.1 shows that the implied unit-level data weights $w_{1},\ldots,w_{n}$ depend only on the treatment indicators, covariates, and error covariance matrix $\bm{\Sigma}$. Therefore, the weights can be obtained as part of the design stage of the study as opposed to the analysis stage. Moreover, the weights can equivalently be obtained by solving a quadratic programming problem.
@@
$\text{subject to }\sum_{i:Z_{i}=1}w_{i}=\sum_{i:Z_{i}=0}w_{i}=1\text{ and }\sum_{i:Z_{i}=1}w_{i}\bm{X}_{i}=\sum_{i:Z_{i}=0}w_{i}\bm{X}_{i}.$ (5)
The solution to this problem is the vector of implied weights of $\widehat{\tau}_{GLS}$,
@@
Proposition 4.2 demonstrates that $\widehat{\tau}_{GLS}$ belongs to the class of minimal dispersion balancing weights estimators, as it arises from solving a weighting problem that minimizes a measure of weight dispersion (4) subject to covariate balance requirements (5). This result builds on the existing literature that shows numerical equivalences between linear regression and balancing weights (Bruns-Smith et al. 2023; Chattopadhyay and Zubizarreta 2023; Kline 2011; Robins et al. 2007) and also relates to the connection between Gaussian process regression, kernel ridge regression, and weighting (Arbour et al. 2021; Kanagawa et al. 2018). To our knowledge, the reformulation of linear regression models with general error structures as simple weighting estimators offers a new perspective into their finite-sample behavior.
@@
The implied weights of GLS sum to one in each treatment arm and exactly balance the means of the measured covariates included in the regression, similar to both weighted least squares (WLS) and augmented inverse probability weighting (AIPW) (Chattopadhyay and Zubizarreta 2023).
@@
A key distinction of the balancing weights problem for GLS, compared to WLS and AIPW, is its objective function $\{\sigma^{2}\sum_{i=1}^{n}w_{i}^{2}+\rho^{2}\sum_{k=1}^{n}\lambda_{k}(\sum_{i:Z_{i}=1}w_{i}v_{ki}-\sum_{i:Z_{i}=0}w_{i}v_{ki})^{2}\}$. This objective simultaneously minimizes the variance of the weights while penalizing differences in the weighted means of a hidden set of covariates — the eigenvectors $\bm{v}_{1},\ldots,\bm{v}_{n}$ of $\bm{S}$ — between the treated and control groups. These eigenvectors correspond to the canonical regressors identified by Hodges and Reich (2010). When $\sigma^{2}$ is small relative to $\rho^{2}\lambda_{1},\ldots,\rho^{2}\lambda_{n}$, the objective function is dominated by the mean differences of eigenvectors with the largest eigenvalues. As a result, GLS approximately balances the means of these eigenvectors across the treated and control groups — that is, $\sum_{i:Z_{i}=1}w_{i}v_{ki}\approx\sum_{i:Z_{i}=0}w_{i}v_{ki}$ for $\lambda_{k}$ large.
@@
Specifically, the minimal weighting problem of Proposition 4.2 admits an equivalent formulation as a convex quadratic program with a soft-balancing constraint on the sum of the mean eigenvector imbalances. One can instead solve $\min_{\bm{w}}\{\sigma^{2}\sum_{i=1}^{n}w_{i}^{2}\}$ subject to $\sum_{i:Z_{i}=1}w_{i}=1,\sum_{i:Z_{i}=0}w_{i}=1\sum_{i:Z_{i}=1}w_{i}\bm{X}_{i}=\sum_{i:Z_{i}=0}w_{i}\bm{X}_{i},$ and $\sum_{k=1}^{n}\lambda_{k}(\sum_{i:Z_{i}=1}w_{i}v_{ki}-\sum_{i:Z_{i}=0}w_{i}v_{ki})^{2}\leq\Delta(\rho^{2})$ for a specific function $\Delta$. The mapping $\rho^{2}\mapsto\Delta(\rho^{2})$ is derived in Supplementary Section 11.2.
@@
The ratio $\sigma^{2}/\rho^{2}$ parametrizes a balance-dispersion trade-off. A larger ratio favors reducing the variance of the weights, whereas a smaller ratio favors balancing the means of eigenvectors with the largest eigenvalues. This result provides a new perspective on the bias-variance tradeoff in random effects and fixed effects models (see Section 11.2.2), and is of independent interest in the more general context of regression analysis.
@@
with respect to $(\bm{\beta},\tau,\bm{\gamma})$ yields $\widehat{\tau}_{\text{ridge}}=\widehat{\tau}_{\text{GLS}}$ (see Supplementary Section 11.3 for details). The ridge penalty on the coefficient for eigenvector $\bm{v}_{k}$ is $\sigma^{2}/(\rho^{2}\lambda_{k})$, so that eigenvectors with the largest eigenvalues are subject to the least shrinkage and effectively enter the linear model as conventional covariates. This perspective also relates to the connection between weighting and kernel ridge regression as discussed in Ben-Michael et al. (2021) and Murray and Feller (2024).
We now interpret this result to explain how spatial regression adjusts for unmeasured spatial confounding. Some intuition can be based on two key observations. First, as discussed above, spatial regression induces approximate mean balance on the eigenvectors of $\bm{S}$ with the largest eigenvalues. Second, these leading eigenvectors correspond to the highest Moran’s I statistics, which is a widely used measure of spatial autocorrelation, where higher values of Moran’s I indicate greater spatial correlation. More formally, the eigenvectors of $\bm{S}$, such as $\bm{v}_{1}=\arg\max_{\bm{x}}\bm{x}^{\top}\bm{S}\bm{x}/\bm{x}^{\top}\bm{x}$, and so forth, maximize the ratio $\bm{x}^{\top}\bm{S}\bm{x}/\bm{x}^{\top}\bm{x}$, which coincides with Moran’s I after centering. Thus, spatial regression implicitly balances the means of a hidden set of unmeasured, spatially autocorrelated covariates. If an unmeasured confounder $\bm{U}$ is highly spatially autocorrelated, it will tend to lie within the span of these eigenvectors, and will therefore be nearly mean-balanced by the spatial regression adjustment. Section 5 formalizes this intuition, providing both a mathematical analysis and visual illustration of the class of unmeasured confounders that spatial regression approximately balances.
@@ 5 Understanding spatial regression contrasts across space
Let $\bm{U}=(U_{1},\ldots,U_{n})$ denote an unmeasured confounder, and let $\overline{U}=\frac{1}{n}\sum_{i=1}^{n}U_{i}$ be its sample mean. In this section, we examine how the bias from $\bm{U}$ depends on its spatial autocorrelation structure.
5.1 Causal identification given $U$
Our target estimand is the average treatment effect on the treated (ATT):
@@ 5.1 Causal identification given $U$
To identify this quantity, we first posit the following general assumptions.
@@
From these assumptions, it follows that
$\tau_{\mathrm{ATT}}=\mathbb{E}\Big{[}\mathbb{E}(Y_{i}|Z_{i}=1,X_{i},U_{i})-\mathbb{E}(Y_{i}|Z_{i}=0,X_{i},U_{i})\ \Big{|}\,Z_{i}=1\Big{]}.$
Now we consider the case when $\bm{U}$ is unobserved. In this setting, adjusting for $\bm{X}$ alone generally yields estimators biased by unmeasured confounding. In what follows, we show that when the outcome model is linear and $\bm{U}$ is spatially autocorrelated, spatial regression methods can control the resulting bias in finite samples.
@@ 5.2 Spatial autocorrelation and unmeasured confounding bias
We now relate the bias from unmeasured confounding by $\bm{U}$ to the degree of its spatial autocorrelation. Moran’s I statistic is a widely used metric for quantifying spatial autocorrelation. Let $\bm{S}$ be an $n\times n$ positive semidefinite spatial weights matrix with eigenvalues $\lambda_{1}\geq\ldots\geq\lambda_{n}\geq 0$. The Moran’s I statistic for $\bm{U}$ relative to $\bm{S}$ is defined as $\mathcal{I}(\bm{U};\bm{S})=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}S_{ij}(U_{i}-\overline{U})(U_{j}-\overline{U})}{\lambda_{1}\sum_{i=1}^{n}(U_{i}-\overline{U})^{2}}$. Although the conventional definition of Moran’s I is $n\lambda_{1}\mathcal{I}(\bm{U};\bm{S})/\sum_{i=1}^{n}\sum_{j=1}^{n}S_{ij}$, the version just presented is normalized to lie between 0 and 1 for ease of interpretation. A value of 1 indicates perfect spatial autocorrelation, while a value of 0 indicates no spatial autocorrelation.
Moran’s I can be interpreted as a scalar summary quantifying how strongly $\bm{U}$ aligns with the eigenvectors of $\bm{S}$ that have the largest eigenvalues. To see this, without loss of generality, assume that $\overline{U}=0$ and $\sum_{i=1}^{n}U_{i}^{2}=1$. Then $\mathcal{I}(\bm{U};S)=\sum_{i=1}^{n}\lambda_{i}(\bm{v}_{i}^{T}\bm{U})^{2}$. In other words, Moran’s I is a weighted sum of the squared coefficients of $\bm{U}$ represented in the eigenbasis of $\bm{S}$, where each coefficient is weighted by the corresponding eigenvalue $\lambda_{i}$. Section 4 suggested that a spatially autocorrelated $\bm{U}$ is largely supported on the leading eigenvectors of $\bm{S}$ with the largest eigenvalues; as a result, $\bm{U}$ is approximately mean-balanced by the weighting procedure of Proposition 4.2. Therefore, spatial regression approximately adjusts for such $\bm{U}$ if the true outcome model is linear. The following proposition formalizes this intuition by relating the bias from unmeasured confounding due to $\bm{U}$ to its Moran’s I statistic.
Proposition 5.1 (Relation between unmeasured confounding bias from $\bm{U}$ and Moran’s I).
@@
$\leq|\gamma|\sqrt{\frac{c_{0}(2\sigma^{2}+\rho^{2}\lambda_{1}+\rho^{2}\lambda_{n})^{2}}{4(\sigma^{2}+\rho^{2}\lambda_{1})(\sigma^{2}+\rho^{2}\lambda_{n})(\sigma^{2}+\rho^{2}\lambda_{1}\mathcal{I}(\bm{U};S))}\sum_{i=1}^{n}(U_{i}-\overline{U})^{2}}$
where $c_{0}=\sigma^{2}\sum_{i=1}^{n}w_{i}^{2}+\rho^{2}\sum_{k=1}^{n}\lambda_{k}(\sum_{i:Z_{i}=1}w_{i}v_{ki}-\sum_{i:Z_{i}=0}w_{i}v_{ki})^{2}$ and $\mathcal{I}(\bm{U};S)$ denotes the Moran’s I statistic of $\bm{U}$ as defined above.
Proposition 5.1 shows that the bias due to unmeasured confounding from $\bm{U}$ tends to decrease as its Moran’s I statistic increases. In other words, spatial regression reduces confounding bias arising from spatially autocorrelated unmeasured confounders. Previous works have arrived at similar results, although through analytical or simulation-based analyses that are less generalizable or interpretable. For example, Paciorek (2010) assume that $\bm{U}$ and $\bm{Z}$ jointly follow a Gaussian process and use the range parameter of the Matérn correlation function to describe spatial autocorrelation, showing that bias decreases as the range of $\bm{U}$ increases. Dupont et al. (2023) quantify spatial smoothness through the eigenvalues of $\bm{S}$ and find that confounding bias is small when $\bm{U}$ aligns with the eigenvectors of $\bm{S}$ corresponding to high eigenvalues. Narcisi et al. (2024) define smoothness in terms of $\bm{U}^{T}\sqrt{\bm{\Sigma}^{-1}}(I-\bm{1}\bm{1}^{T}/n)\sqrt{\bm{\Sigma}^{-1}}\bm{U}$, and show that bias decreases as this measure increases. In contrast, Proposition 5.1 avoids assumptions regarding the distribution or structure of $\bm{U}$ and $\bm{S}$, and makes use of Moran’s I, which is a widely used measure of spatial autocorrelation.

### v2

@@ 3.3 Gaussian process models
We apply the GP model to the Superfund dataset by setting $\rho^{2}=10$ and $\sigma^{2}=1$ and using a Matérn correlation with hyperparameters $\kappa=10,\phi=500/(2\sqrt{10})$ and the distance $d_{ij}$ between sites $i$ and $j$ is measured in meters. Specifically, $K_{ij}=2^{1-\kappa}(\Gamma(\kappa))^{-1}(d_{ij}/\phi)^{\kappa}B_{\kappa}(d_{ij}/\phi)$ where $B_{\kappa}$ is the modified Bessel function of the second kind of order $\kappa$.
@@ 4 An encompassing weighting framework
This section presents our main results. First, we establish that the GLS estimator $\widehat{\tau}_{GLS}$ in Equation 2 acts as a weighting estimator. Second, we demonstrate that it is specifically a minimal dispersion approximately balancing weights estimator: the weights minimize a measure of weight dispersion while achieving exact balance on the means of measured covariates and approximate balance on the means of a latent set of covariates. This property elucidates the exact nature of the adjustment induced by GLS, clarifying how it approximates key features of a hypothetical experiment and distinguishing it from other forms of regression adjustment.
While the following propositions apply to any error covariance matrix $\bm{\Sigma}$, we interpret them here in the context of spatial regression, where $\bm{\Sigma}=\sigma^{2}\bm{I}_{n}+\rho^{2}\bm{S}$ and $\bm{S}$ encodes spatial autocorrelation. The following proposition shows that $\widehat{\tau}_{GLS}$ admits a weighting representation, as it can be re-expressed as a difference of weighted means of the treated and control outcomes.
@@
For any positive semidefinite covariance matrix $\Sigma$, the GLS estimator of $\tau$ in Equation 2 can be expressed as
@@
with weights of $\bm{w}=(w_{1},\ldots,w_{n})$ that admit the following closed-form representation
@@
Proposition 4.1 shows that the implied unit-level data weights $w_{1},\ldots,w_{n}$ depend only on the treatment indicators, covariates, and error covariance matrix $\bm{\Sigma}$. Therefore, the weights can be obtained as part of the design stage of the study as opposed to analysis stage. Moreover, the weights can equivalently be obtained by solving a quadratic programming problem.
@@
$\text{subject to }\sum_{i:Z_{i}=1}w_{i}\bm{X}_{i}=\sum_{i:Z_{i}=0}w_{i}\bm{X}_{i},\sum_{i:Z_{i}=1}w_{i}=\sum_{i:Z_{i}=0}w_{i}=1.$ (5)
The solution to this problem is the implied weights of $\widehat{\tau}_{GLS}$,
@@
Proposition 4.2 demonstrates that $\widehat{\tau}_{GLS}$ belongs to the class of minimal dispersion approximately balancing weights estimators, as it arises from solving a weighting problem that minimizes a measure of weight dispersion (4) subject to covariate balance requirements (5). An extension to a more general $\bm{\Sigma}$, allowing any positive semidefinite structure, is given in Corollary A.1 in the Supplementary Material. This result builds on the existing literature that shows numerical equivalences between linear regression and balancing weights (Bruns-Smith et al. 2023; Chattopadhyay and Zubizarreta 2023; Robins et al. 2007) and also relates to the connection between Gaussian process regression, kernel ridge regression, and weighting (Arbour et al. 2021). To our knowledge, the reformulation of linear regression models with general error structures as weighting estimators offers a new perspective into their finite-sample behavior.
@@
The implied weights of GLS sum to one in each treatment arm and exactly balance the means of the measured covariates included in the regression, similar to both weighted least squares and augmented inverse probability weighting (Chattopadhyay and Zubizarreta 2023).
@@
A key distinction of the balancing weights problem for GLS, compared to weighted least squares and augmented inverse probability weighting, is its objective function $\{\sigma^{2}\sum_{i=1}^{n}w_{i}^{2}+\rho^{2}\sum_{k=1}^{n}\lambda_{k}(\sum_{i:Z_{i}=1}w_{i}v_{ki}-\sum_{i:Z_{i}=0}w_{i}v_{ki})^{2}\}$. This objective simultaneously minimizes the variance of the weights while penalizing differences in the weighted means of a latent set of covariates—the eigenvectors $\bm{v}_{1},\ldots,\bm{v}_{n}$ of $\bm{S}$—between the treated and control groups. These eigenvectors correspond to the canonical regressors identified by Hodges and Reich (2010). When $\sigma^{2}$ is small relative to $\rho^{2}\lambda_{1},\ldots,\rho^{2}\lambda_{n}$, the objective function is dominated by the mean differences of eigenvectors with the largest eigenvalues. As a result, GLS approximately balances the means of these eigenvectors across the treated and control groups—that is, $\sum_{i:Z_{i}=1}w_{i}v_{ki}\approx\sum_{i:Z_{i}=0}w_{i}v_{ki}$ for $\lambda_{k}$ large.
@@
Specifically, the weighting problem of Proposition 4.2 admits an equivalent formulation as a convex quadratic program with a soft‐balancing constraint on the sum of the mean eigenvector imbalances. In particular, one can instead solve $\min_{\bm{w}}\{\sigma^{2}\sum_{i=1}^{n}w_{i}^{2}\}$ subject to $\sum_{i:Z_{i}=1}w_{i}=1,\sum_{i:Z_{i}=0}w_{i}=1\sum_{i:Z_{i}=1}w_{i}\bm{X}_{i}=\sum_{i:Z_{i}=0}w_{i}\bm{X}_{i},$ and $\sum_{k=1}^{n}\lambda_{k}(\sum_{i:Z_{i}=1}w_{i}v_{ki}-\sum_{i:Z_{i}=0}w_{i}v_{ki})^{2}\leq\Delta(\rho^{2})$ for a specific function $\Delta$. The mapping $\rho^{2}\mapsto\Delta(\rho^{2})$ is derived in Supplementary Section B.
@@
The ratio $\sigma^{2}/\rho^{2}$ thus parametrizes a balance-dispersion trade-off. A larger ratio favors reducing the variance of the weights, whereas a smaller ratio favors balancing the means of eigenvectors with the largest eigenvalues. This result provides a new perspective on the bias-variance tradeoff in random effects and fixed effects models (see Supplementary Section B), and is of independent interest in the more general context of regression analysis.
@@
with respect to $(\bm{\beta},\tau,\bm{\gamma})$ yields $\widehat{\tau}_{\text{ridge}}=\widehat{\tau}_{\text{GLS}}$ (see Supplementary Section C for details). The ridge penalty on the coefficient for eigenvector $\bm{v}_{k}$ is $\sigma^{2}/(\rho^{2}\lambda_{k})$, so that eigenvectors with the largest eigenvalues are subject to the least shrinkage and effectively enter the linear model as covariates. This perspective also relates to the connection between weighting and kernel ridge regression as discussed in Ben-Michael et al. (2021) and Murray and Feller (2024).
We now interpret this result to explain how spatial regression adjusts for confounding. Some intuition can be based on two key observations. First, as discussed above, spatial regression induces approximate mean balance on the eigenvectors of $\bm{S}$ with the largest eigenvalues. Second, these leading eigenvectors correspond to the highest Moran’s I statistics, which is a widely used measure of spatial autocorrelation, where higher values of Moran’s I indicate greater spatial correlation. More formally, the first eigenvector $\bm{v}_{1}$ of $\bm{S}$ maximizes the ratio $\bm{x}^{\top}\bm{S}\bm{x}/\bm{x}^{\top}\bm{x}$, which coincides with Moran’s I after centering; $\bm{v}_{2}$ maximizes the ratio in the subspace orthogonal to $\bm{v}_{1}$, and so forth. Thus, spatial regression implicitly balances the means of a latent set of unmeasured, spatially autocorrelated covariates. If an unmeasured confounder $\bm{U}$ is highly spatially autocorrelated, it will tend to lie within the span of these eigenvectors, and will therefore be nearly mean-balanced by the spatial regression adjustment. Section 5 formalizes this intuition, providing both a mathematical analysis and visual illustration of the class of unmeasured confounders that spatial regression approximately balances.
@@ 5 Understanding spatial regression contrasts across space
Let $\bm{U}=(U_{1},\ldots,U_{n})$ denote an unmeasured confounder with sample mean $\overline{U}=\sum_{i=1}^{n}U_{i}/n$. In this section, we examine how the bias induced by $\bm{U}$ depends on its spatial autocorrelation structure. Our target estimand is the average treatment effect on the treated (ATT):
@@
5.1 Causal identification
To identify this quantity, we first posit the following assumptions.
@@ 5.1 Causal identification
Under these assumptions, $\tau_{\mathrm{ATT}}=\mathbb{E}\Big[\mathbb{E}(Y_{i}|Z_{i}=1,X_{i},U_{i})-\mathbb{E}(Y_{i}|Z_{i}=0,X_{i},U_{i})\ \Big|\,Z_{i}=1\Big].$ However, $\bm{U}$ is unobserved and an analysis adjusting only for $\bm{X}$ will be subject to bias. In what follows, we show that when the outcome model is linear and $\bm{U}$ has a specific form of spatial autocorrelation, spatial regression methods can limit the magnitude of the resulting bias in finite samples.
@@ 5.2 Spatial autocorrelation and unmeasured confounding bias
We now relate the bias due to the unmeasured confounder $\bm{U}$ to the strength of its spatial autocorrelation. Moran’s I statistic is a widely used metric for quantifying spatial autocorrelation. Let $\bm{S}$ be an $n\times n$ positive semidefinite spatial weights matrix with eigenvalues $\lambda_{1}\geq\ldots\geq\lambda_{n}\geq 0$. The Moran’s I statistic for $\bm{U}$ relative to $\bm{S}$ is defined as $\mathcal{I}(\bm{U};\bm{S})=\sum_{i=1}^{n}\sum_{j=1}^{n}S_{ij}(U_{i}-\overline{U})(U_{j}-\overline{U})/(\lambda_{1}\sum_{i=1}^{n}(U_{i}-\overline{U})^{2})$. Although the conventional definition of Moran’s I is $n\lambda_{1}\mathcal{I}(\bm{U};\bm{S})/\sum_{i=1}^{n}\sum_{j=1}^{n}S_{ij}$, the version just presented is normalized to lie between 0 and 1 for ease of interpretation. A value of 1 indicates perfect spatial autocorrelation, while a value of 0 indicates no spatial autocorrelation.
Moran’s I can be interpreted as a scalar summary quantifying how strongly $\bm{U}$ aligns with the eigenvectors of $\bm{S}$ that have the largest eigenvalues. To see this, without loss of generality, assume that $\overline{U}=0$ and $\sum_{i=1}^{n}U_{i}^{2}=1$. Then $\mathcal{I}(\bm{U};\bm{S})=\sum_{i=1}^{n}\lambda_{i}(\bm{v}_{i}^{T}\bm{U})^{2}$. In other words, Moran’s I is a weighted sum of the squared coefficients of $\bm{U}$ represented in the eigenbasis of $\bm{S}$, where each coefficient is weighted by the corresponding eigenvalue $\lambda_{i}$. Section 4 showed that a spatially autocorrelated $\bm{U}$ is largely supported on the leading eigenvectors of $\bm{S}$ with the largest eigenvalues; as a result, $\bm{U}$ is approximately mean-balanced by the weighting procedure of Proposition 4.2. Therefore, spatial regression approximately adjusts for such $\bm{U}$ if the true outcome model is linear. The following proposition formalizes this intuition by relating the bias from unmeasured confounding due to $\bm{U}$ to its Moran’s I statistic.
Proposition 5.1 (Relation between Moran’s I and unmeasured confounding bias from $\bm{U}$).
@@
$\leq|\gamma|\sqrt{\frac{c_{0}(2\sigma^{2}+\rho^{2}\lambda_{1}+\rho^{2}\lambda_{n})^{2}}{4(\sigma^{2}+\rho^{2}\lambda_{1})(\sigma^{2}+\rho^{2}\lambda_{n})(\sigma^{2}+\rho^{2}\lambda_{1}\mathcal{I}(\bm{U};\bm{S}))}\sum_{i=1}^{n}(U_{i}-\overline{U})^{2}}$
where $c_{0}=\sigma^{2}\sum_{i=1}^{n}w_{i}^{2}+\rho^{2}\sum_{k=1}^{n}\lambda_{k}(\sum_{i:Z_{i}=1}w_{i}v_{ki}-\sum_{i:Z_{i}=0}w_{i}v_{ki})^{2}$ and $\mathcal{I}(\bm{U};\bm{S})$ denotes the Moran’s I statistic of $\bm{U}$ as defined above.
Proposition 5.1 shows that the bias due to unmeasured confounding from $\bm{U}$ tends to decrease as its Moran’s I statistic increases. In other words, spatial regression reduces confounding bias arising from spatially autocorrelated unmeasured confounders. Previous works have arrived at similar results, although through analytical or simulation-based analyses that are less generalizable or interpretable. For example, Paciorek (2010) assume that $\bm{U}$ and $\bm{Z}$ jointly follow a Gaussian process and use the range parameter of the Matérn correlation function to describe spatial autocorrelation, showing that bias decreases as the range of $\bm{U}$ increases. Dupont et al. (2023) quantify spatial smoothness through the eigenvalues of $\bm{S}$ and find that confounding bias is small when $\bm{U}$ aligns with the eigenvectors of $\bm{S}$ corresponding to high eigenvalues. Narcisi et al. (2024) define smoothness in terms of $\bm{U}^{T}\sqrt{\bm{\Sigma}^{-1}}(I-\bm{1}\bm{1}^{T}/n)\sqrt{\bm{\Sigma}^{-1}}\bm{U}$, and show that bias decreases as this measure increases. By contrast, Proposition 5.1 avoids assumptions regarding the distribution or structure of $\bm{U}$ and $\bm{S}$, and makes use of Moran’s I, which is a widely used measure of spatial autocorrelation.